A regression was run to determine if there is a relationship between hours of TV watched per day (x) and number of situps a person can do (y). The results of the regression were: y=ax+b a=-1.179 b=22.006 r2=0.595984 r=-0.772 Use this to predict the number of situps a person who watches 5.5 hours of TV can do (to one decimal place)
Answer #1

hours of TV watched per day (x)

number of situps a person can do (y).

The results of the regression were:

y=ax+b

a=-1.179 b=22.006

r2=0.595984 . r=-0.772

Given :- X=5.5

Y^ = -1.179 * X + 22.006

Y^ = -1.179 * 5.5 + 22.006 = 15.5